### v1

# Preference-free Alignment Learning with Regularized Relevance Reward

Abstract
Learning from human preference has been considered key to aligning Large Language Models (LLMs) with human values. However, contrary to popular belief, our preliminary study reveals that reward models trained on human preference datasets tend to give higher scores to long off-topic responses than short on-topic ones. Motivated by this observation, we explore a preference-free approach utilizing ‘relevance’ as a key objective for alignment. On our first attempt, we find that the relevance score obtained by a retriever alone is vulnerable to reward hacking, i.e., overoptimizing to undesired shortcuts, when we utilize the score as a reward for reinforcement learning. To mitigate it, we integrate effective inductive biases into the vanilla relevance to regularize each other, resulting in a mixture of reward functions: Regularized Relevance Reward (R). R significantly improves performance on preference benchmarks by providing a robust reward signal. Notably, R does not require any human preference datasets (i.e., preference-free), outperforming open-source reward models in improving human preference. Our analysis demonstrates that R has advantages in elevating human preference while minimizing its side effects. Finally, we show the generalizability of R, consistently improving instruction-tuned models in various backbones and sizes without additional dataset cost 111Our code is available at github.com/naver-ai/RRR..
1 Introduction
To align large language models (LLM) with human values like helpfulness, human preference-based learning has been studied (Ouyang et al., 2022; Bai et al., 2022a, b; Rafailov et al., 2023). Typically, the preference is represented as a pair of responses where one is considered the chosen option, and the other is a rejected option for a given input under a Bradley-Terry model (Bradley & Terry, 1952). As illustrated in the upper side of Figure 1, Reinforcement Learning from Human Feedback (RLHF) is a dominant approach to exploiting pairwise feedback. The pairwise feedback is used to train a proxy reward model (RM), and a policy model is optimized over the reward signal from the RM using RL (Schulman et al., 2017; Ziegler et al., 2020). Also, there have been introduced alternative approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Sequence Likelihood Calibration (SLiC) (Zhao et al., 2023), which directly optimize preference pairs while not requiring an explicit RM.
On the other hand, a line of research has studied the innate limitations of human preference (feedback). Xu et al. (2023) reveal prior knowledge affects the preference judgment in that crowd workers often choose the preferred response considering surface-level properties such as conciseness and specificity, while experts focus on more essential properties like factuality and completeness. Similarly, Hosking et al. (2023) further categorize the attributions of the preference judgments and confirm that human judgments could be biased by stylistic features like assertiveness. Similarly, during our preliminary study, we find that most open-source RMs (Köpf et al., 2023; Ethayarajh et al., 2023a; Touvron et al., 2023; Cui et al., 2023), trained with human preference, struggle with discriminating responses that are more related or on-topic with the given input, while an unsupervised retriever (Izacard et al., 2021) perfectly distinguishes them (Figure 2 shows an example.). This observation makes us wonder whether relevance can be used for the LLM alignment without explicit human preference, i.e., directly using it as a reward signal.
In this work, we present a preference-free approach focusing on relevance, motivated by our observation. We define ‘Relevance’ as the degree to which an assistant system’s responses are appropriate and directly connected to the specific needs or requests presented by the users. As shown in Figure 1, our approach does not require any preference dataset to align LLMs. Instead, we propose a mixture of reward functions: Regularized Relevance Reward (R). We show that when regularized appropriately, the relevance between the input and output could be a reliable learning signal. Specifically, we integrate a set of effective inductive biases such as length incentive, repetition penalty, and distinction by query type into the relevance score to suppress the risk of reward hacking, which is a phenomenon to exploit undesired shortcuts in the reward signal (Skalse et al., 2022; Pang et al., 2022). We use R as a reward function to align LLMs within Proximal Policy Optimization (PPO) (Schulman et al., 2017; Ziegler et al., 2020).
Experimental results show that our method significantly improves human preference over LLMs even without explicit human or AI feedback, outperforming state-of-the-art open-source RMs (Köpf et al., 2023; Ethayarajh et al., 2023a; Touvron et al., 2023; Cui et al., 2023). Also, we offer an in-depth analysis to compare the resulting response distribution between models trained with preference-based RM and our R, focusing on relevance. We find that all PPO models tend to generate irrelevant sentences as the response gets longer, but the model trained with R produces more relevant responses. Further ablation study demonstrates that R has advantages in improving preference while minimizing alignment tax (Ouyang et al., 2022), a phenomenon where the increased preference accompanies the degraded performance on other NLP tasks. Finally, R consistently improves aligned behaviors of strong instruction-following (SFT) models, Vicuna and Mistral (Chiang et al., 2023; Jiang et al., 2023; Tunstall et al., 2023), using only input-output pairs, i.e., preference-free.
2 Motivation and Preliminary Study
In alignment learning, human preference could be divided into several facets (Ye et al., 2023; Kim et al., 2023b), e.g., fluency, accuracy, completeness, relevance of the responses, etc. Among them, Relevance mainly refers to the degree to which the responses of an assistant system are appropriate and directly connected to the specific needs or requests presented by the users. It emphasizes the system’s ability to understand and appropriately address the user’s intent and context. Our study focuses on the response’s relevance with the provided query from the users, starting by asking the following two questions: (1) Can open-source RMs capture relevance? and (2) Can relevance-only be used for alignment?
2.1 Can open-source RMs capture relevance?
First, we study how open-source reward models (RM) capture relevance between query and response behind an overall preference. A common belief is that the reward model is capturing relevance well, but it turns out that (1) it does not, and (2) a simple relevance model like Contriever (Izacard et al., 2021) does it much better out of the box. For this study, we devise an adversarial test set, Synthetic Relevance (SynRel), containing 530 (query, relevant response, and irrelevant response) triplets, as illustrated in Figure 2. Specifically, we randomly sample 530 entities whose popularity is less than 1k in PopQA (Mallen et al., 2023) 222The PopQA dataset contains the quantified popularity of each entity, e.g., monthly page-view of Wikipedia.. The query for each instance is formed using the template “Please tell me about {entity}”, and the positive (chosen) response is created by compiling the ground truth property of the entity, e.g., “{entity} is {property}”. The negative (rejected) response is generated by prompting another negative entity to GPT-3.5-turbo (OpenAI, 2023).
We compare four RMs, OASST-RM 333OpenAssistant/reward-model-deberta-v3-large-v2 (Köpf et al., 2023), SteamSHP-XL (Ethayarajh et al., 2023a), LLaMA-2-7B-RM (Touvron et al., 2023), which is our implementation trained on Anthropic-HH (Bai et al., 2022a) and SHP (Ethayarajh et al., 2023a), and UltraRM-13B (Cui et al., 2023), on the SynRel test set. Also, we include Contriever (Izacard et al., 2021), originally used as a retriever, as a representative model trained to capture relevance. We report preference accuracy for selecting the chosen response for a given query in Table 1. We find that the test set of SynRel is challenging for most open-source RMs, even though they perform well in other preference datasets, Anthropic-HH and SHP. In other words, the reward models trained on overall human preference cannot capture the trivial relevance problems that a far smaller retriever model can perfectly distinguish 444The parameter size of UltraRM (13B) is about 118x larger than that of Contriever (110M)..
RM Overall Preference Relevance Anthropic-HH SHP SynRel Random 50.0 50.0 50.0 OASST-RM 67.6 53.9 52.6 SteamSHP-XL 68.5 74.8 58.9 LLaMA2-7B-RM 69.6 72.3 45.1 UltraRM-13B 70.8 73.4 59.8 Contriever 54.0 56.0 100
2.2 Can relevance-only be used for alignment?
From the observation of the previous subsection, we wonder if relevance could be used as a meaningful learning signal for LLM alignment. We could think of two ways to implement this. The first is training a canonical RM on relevance dataset (Wu et al., 2023), and the second is using an off-the-shelf retriever directly for the RM. Here, we choose the second option using Contriever denoted with for simplicity. Let be a query and be a rollout response in the exploration stage of Proximal Policy Optimization (PPO) (Schulman et al., 2017; Ziegler et al., 2020). We compute the relevance score between the query and sampled response: .
We use the Anthropic-HH (Bai et al., 2022a) for the experiments. We start by training supervised policy, i.e., supervised fine-tuning (SFT), using chosen responses in the dataset. Then, we perform PPO training upon the SFT model against relevance score . Please see Appendix C for more implementation details. Finally, We conduct GPT-4 evaluation (OpenAI, 2023) to inspect whether there are preference improvements after the RL training compared to the SFT model. Specifically, we compare the SFT and PPO models with a pair-wise judgment of Vicuna-Bench (Chiang et al., 2023). We report the adjusted win rate computed by (# Win + 0.5 # Tie) / # All) comparing the PPO models with the SFT baseline.
Model Vicuna-Bench % Adjusted Win vs. SFT Relevance score Avg # Tokens SFT 50.0 4.94 173 + PPO w. OASST-RM 61.3 6.49 174 w. UltraRM-13B 73.8 6.34 227 w. 20.0 8.06 69
In Table 2, we find that solely using can not improve the preference and even worsens the model. Contrary to other RMs, the relevance reward significantly drops the adjusted win rate in the Vicuna-Bench. We further analyze the problem by measuring query-response relevance score and response length. For this, we employ another open-source relevance model trained as a reranker 555huggingface.co/cross-encoder/ms-marco-MiniLM-L-12-v2 to compute the relevance score. The PPO model trained using the demonstrates the highest relevance score; however, its responses are notably brief in length. We speculate using only for the reward is vulnerable to reward hacking (Skalse et al., 2022; Pang et al., 2022). For example, the best strategy to maximize would be just copying the question for the response. In fact, about 40% of the generated responses are merely copies of the given questions. Also, it reminds us of the trade-off relationship between relevance and the short response length discussed in Wu et al. (2023).
3 R: Regularized Relevance Reward
To overcome the easily hackable nature of the vanilla relevance reward , we propose a Regularized Relevance Reward (R), the mixture of reward functions, including the that regularize each other. Note that this is just one way to regularize the vanilla relevance reward with a strong inductive bias, and one could explore better ways. We will demonstrate each function in this Section.
Length Incentive (LI)
First, we introduce Length Incentive (LI) to mitigate the ’s tendency to shorten the length of responses. It scores the reward by calculating . Intuitively, this reward function promotes the longer generation regardless of the given query . It is similar to LPPO suggested by concurrent work (Singhal et al., 2023). However, we do not normalize the reward with the target length. We find that combining LI with the is somewhat effective in improving preference. However, in this setting, we encounter another kind of reward hacking, undesired repetitions in the generations. For example, the best strategy to maximize the is just repeating the .
Repetition Penalty (RP)
Hence, we additionally introduce the Repetition penalty (RP) to regularize the undesired outputs. It checks the unique trigram ratio of , i.e., . Integrating it with the reward functions efficiently prevents redundancy and encourages proper length of outputs. The intuition behind the mixture of reward functions is aiming for high relevance between query and generated response while promoting long response length and less repetitions. However, the Length Incentive could be problematic when the question requires constrained or factual responses. In these cases, consistency of response might be much more important than the long and diverse responses, as discussed in Touvron et al. (2023); Song et al. (2023b).
Reference Answer Relevance ()
We apply the different relevance rewards according to query type to handle such cases. Specifically, we define two types of queries, Open-ended (OE) and Closed-ended (CE), according to whether they require creative and open-ended or consistent and constrained responses, similar to Song et al. (2023b). Then, we include reference answer relevance , where the is the ground-truth reference response of given query.
Finally, we present our final reward function design, Regularized Relevance Reward (R):
, where the indicates the pre-identified query types and the indicates a linear interpolation function mapping the range of to the range of for stable RL training. The mixture of reward functions regularizes each other, making it robust to reward hacking. Besides, It only requires an SFT dataset consisting of and pairs, eliminating the need for an extensive feedback dataset to train an additional reward model.
Query-Type Classification
Deciding whether the given query requires open-ended generation may not be obvious because of its subjectivity in interpreting the range of ‘open-ended’. Nevertheless, we define the meaning of open-ended so that the responses to the query can be relatively anything and creative. For example, “How can I make a good first impression?” could be an open-ended query. On the other side, the closed-ended query indicates the corresponding responses should be consistent and constrained as a closed form. “How many movies are in the Fast and Furious franchise?” would be one of the closed-ended queries. Considering resource constraints, we automatically annotate the query type (OE or CE) used in R by prompting it to GPT-4 (OpenAI, 2023). We find that about 83% of decisions by GPT-4 are agreed upon when the authors manually validate 100 samples in a preliminary stage. Also, when we distill the classification task into in-house LLM, LLaMA-2-7B (Touvron et al., 2023), we achieve competitive accuracy in the test set 666When we do not specify otherwise, is derived from GPT-4 in our experiments.. Please see Appendix B for more details of the query-type classification.
PPO training against R
We use the resulting mixture of reward functions, R, to further optimize an SFT model with reinforcement learning. Consequently, the following PPO objective induces the model to maximize R,
while online policy and the fixed reference (SFT) policy do not diverge too much by regularizing it with KL constraints (Schulman et al., 2017; Ziegler et al., 2020), where the is train dataset and the is the coefficient to control the KL penalty.
Model # Feedback to train RM Vicuna-Bench % Adjusted Win vs. SFT AlpacaEval % Win vs. Text-Davinci-003 MT-Bench 1st / 2nd / Avg ToxiGen () SFT (Anthropic-HH) - 50.0 15.9 4.6 / 3.4 / 3.9 2.6 + PPO w. SteamSHP-XL 477k 59.4 19.6 4.2 / 3.4 / 3.8 6.2 w. OASST 393k 61.3 19.1 4.7 / 3.1 / 3.9 5.4 w. LLaMA2-7B-RM 509k 62.5 19.8 4.6 / 3.5 / 4.0 5.6 w. UltraRM-13B 256k 73.8 28.0 4.9 / 3.6 / 4.2 4.7 w. R (Ours) 0 76.9 23.4 5.0 / 3.7 / 4.4 3.9
4 Experiments
We demonstrate the efficacy of R as a reliable reward for alignment (Section 4.2), further advantages of R with ablation study (4.3), and its generalizability as a preference-free alignment approach (Section 4.4).
4.1 Common Experimental Setup
We mainly conduct experiments based on two alignment datasets, Anthropic-HH (Bai et al., 2022a) and AlpacaFarm (Dubois et al., 2023). Also, we train three instruction-tuned models using R with corresponding SFT datasets. We perform PPO (Ziegler et al., 2020) training with rewards from R utilizing only () pairs, i.e., no preference triplets. As in Section 2, we employ Contriever (Izacard et al., 2021) for scoring and 777huggingface.co/facebook/contriever. More implementation details are in Appendix C. We evaluate the resulting models on three GPT-4-based benchmarks for alignment evaluation: Vicuna-Bench (Chiang et al., 2023), MT-Bench (Zheng et al., 2023), and AlpacaEval (Li et al., 2023), and one classifier-based safety benchmark, ToxiGen (Hartvigsen et al., 2022). In addition, we include SuperNI (Wang et al., 2022) and FactScore (Min et al., 2023) to measure alignment tax. More evaluation details are in Appendix D.
4.2 Can R improve human preference?
We compare R with open-source RMs in improving human preference, e.g., producing more helpful and harmless responses, based on Anthropic-HH dataset (Bai et al., 2022a).
Baseline RMs
We include the open-source RMs studied in Section 2 as baseline RMs for PPO training (Ziegler et al., 2020). OpenAssistant (OASST) is trained with four public feedback datasets, WebGPT (Nakano et al., 2021), OpenAI Summary (Stiennon et al., 2020), Anthropic-HH (Bai et al., 2022a), and SyntheticGPT 888Dahoas/synthetic-instruct-gptj-pairwise, based on DeBERTa-V3-Large (He et al., 2021). SteamSHP-XL is another open-source RM trained with SHP (Ethayarajh et al., 2023a) and Helpfulness subset of the Anthropic-HH. Unlike other feedback datasets, the SHP consists of human-generated responses from an online community, Reddit. It is based on FLAN-T5-XL (Chung et al., 2022). Also, we train RM baseline based on LLaMA-2-7B (Touvron et al., 2023) using Anthropic-HH and SHP, LLaMA-2-7B-RM. We include UltraRM-13B, the most capable open-source RM trained on UltraFeedback (Cui et al., 2023). The UltraFeedback is a 64k AI feedback dataset constructed by GPT-4 (OpenAI, 2023). Also, it is further fine-tuned with the public feedback datasets, Anthropic-HH, SHP, and OpenAI Summary. Based on the same experimental setup, we perform PPO training against the reward score by each RM or R.
R is effective for alignment
In Table 3, we can find that our R significantly improves the model’s preference compared to the SFT baseline on all the preference benchmarks, Vicuna, Alpaca, and MT. Surprisingly, R surpasses all the open-source RMs in the Vicuna, MT-Bench, and ToxiGen even though R is a preference-free approach, i.e., zero feedback dataset 999In Appendix F, R outperforms other preference-based learning approaches, including DPO (Rafailov et al., 2023).. As we use only the helpful subset for the PPO, all the PPO models increase their toxicity compared to the SFT baseline. However, our model is found to be the safest among the PPO models, indicating our improvements on the preference benchmarks are achieved while retaining safety. We presume our lower performance in AlpacaEval compared to UltraRM is because AlpacaEval includes neither the reference answer nor the bi-positional judgments, unlike Vicuna-Bench and MT-Bench. On the other hand, we find that PPO training with R increases reward score by UltraRM as well, as shown in Figure 11 of Appendix E.1.
Model AlpacaEval % Win / % Rel SuperNI ROUGE-L FactScore Score / # Facts SFT (AlpacaFarm) 23.2 / 88.2 34.9 38.4 / 19.9 + PPO w. UltraRM 34.1 / 85.1 28.8 33.6 / 25.8 w. R 37.2 / 86.4 34.9 36.8 / 21.4 w. random 34.8 / 82.6 28.4 36.7 / 22.1 w. learned 38.4 / 85.9 31.2 36.4 / 23.1 w. R () 41.5 / 84.8 22.6 33.3 / 42.4 w. 41.6 / 81.3 13.1 32.4 / 39.6 w. 24.1 / 64.5 7.4 33.9 / 56.6
R increases preference while keeping query relevance
We further investigate the results of AlpacaEval, focusing on the difference between response distributions. In particular, we report Relevant Sentence Ratio (% Rel) of models’ responses to compare their distribution related to each component of R. We measure the % Rel by asking whether each sentence in the generated response is relevant to the given query to GPT-4 (OpenAI, 2023). We request GPT-4 to refer to each sentence first and then judge its query relevance as either ‘Relevant’ or ‘Irrelevant’. That is, it indicates sentence-level query relevance of the responses. The regarding prompt for this evaluation is found in Appendix D.2. We measure % Rel of the models based on the AlpacaEval set. As a result, R results in a more relevant response (85.1%) than all the open-source RMs, including UltraRM (83.1%), trained with AI and human preferences 101010More detailed results are found in Table 8.. We visualize their differences in sentence-level relevance in Figure 3. It exhibits scatter plots of the % Rel according to sentence order. We find all the models tend to show a high relevance at the starting point of each response. However, both PPO models (trained with either UltraRM or R) produce irrelevant sentences as the response gets longer. When comparing the two PPO models, the model trained with UltraRM tends to generate longer off-topic content than R. Also, we find that an ensemble of rewards by R and UltraRM further improves the relevance in Appendix E.2.
4.3 Is the reward design of R effective?
We validate R’s reward design choice by ablating each component one by one from the R. We conduct these experiments based on another alignment dataset, AlpacaFarm (Dubois et al., 2023). We further evaluate models on SuperNI (Wang et al., 2022) and Factscore (Min et al., 2023), containing many closed-ended (CE) queries. We intend to capture alignment tax (Ouyang et al., 2022), the trade-offs for the increased preference from the evaluations.
Backbone Model Alignment Method Feedback (Type / #) AlpacaEval % Win / % Rel MT-Bench 1st / 2nd / Avg ToxiGen () SuperNI ROUGE-L Llama-2-7B Llama-2-7B-Chat PPO Human / 1.4M 71.37 / 86.7 6.13 / 5.81 / 5.97 0.0 7.7 Vicuna-7B-v1.5 SFT - / 0 73.72 / 91.5 6.76 / 5.38 / 6.07 0.3 34.3 + R (Ours) PPO / LoRA - / 0 79.13 / 84.4 7.02 / 5.86 / 6.44 0.1 25.8 Mistral-7B Zephyr-beta DPO AI / 62k 87.80 / 92.2 7.65 / 6.75 / 7.20 0.5 13.5 Mistral-H4-SFT SFT - / 0 75.34 / 91.6 6.88 / 6.11 / 6.49 1.5 22.4 + R (Ours) PPO / LoRA - / 0 77.33 / 90.8 7.20 / 6.06 / 6.63 1.4 15.7 Llama-2-13B Llama-2-13B-Chat PPO Human / 1.4M 78.11 / 95.5 6.96 / 6.24 / 6.60 0.0 7.1 Vicuna-13B-v1.5 SFT - / 0 80.16 / 92.1 6.75 / 6.39 / 6.57 0.2 36.0 + R (Ours) PPO / LoRA - / 0 85.14 / 92.6 7.26 / 6.62 / 6.94 0.1 17.4
R pays less alignment tax
In Table 4, we can find R has an advantage in alignment tax (Ouyang et al., 2022), a phenomenon in which the improved preference accompanies degraded performances on other NLP tasks. For example, R reliably improves the preference (% Win) on the AlpacaEval while keeping high relevance (% Rel), zero-shot instruction-following ability (SuperNI), and factuality (FactScore) simultaneously. However, UltraRM significantly sacrifices the performances in the SuperNI and FactScore compared to the SFT baseline, resulting in lower % Win and % Rel than ours. We compare results according to the choice of relevance model in Appendix G.
Relevance contributes to minimize alignment tax
Moreover, we can observe a clear pattern when excluding the relevance scores from the reward. If we do not use according to query type, i.e., R, it shows even higher preference scores with decreased relevance. However, the achievement accompanies significant performance drops on the SuperNI and FactScore by about 12.3% and 3.5% points, respectively. If the is discarded (), the performance of the relevant sentence ratio and ROUGE-L score of SuperNI drop more drastically. Finally, using only the response length as a reward () improves the preference slightly compared to SFT, as reported in Singhal et al. (2023), but it significantly worsens the % Rel and SuperNI score in return.
Effect of branching by query-type
If we use random query type, random , for the R instead of the predicted one by GPT-4, its win rate and relevant sentence ratio are degraded on the AlpacaEval. Also, its instruction-following ability is sacrificed, indicating that reward differentiation based on query types is effective. On the other hand, we show that R’s query type classification by GPT-4 can be effectively replaced with in-house LLM, LLaMA-2-7B (Touvron et al., 2023), i.e., learned 111111The training details of the in-house LLM are in Appendix B.2.. It achieves an even better win rate than the original R with competitive performances for SuperNI and FactScore. Additionally, we compare the models’ diversity and consistency according to the query type in Figure 4. We report the Self-BLEU (Zhu et al., 2018) by sampling 10 responses for each input in the AlpacaEval 121212We perform nucleus sampling (Holtzman et al., 2019).. A high Self-BLEU score indicates a consistent response, whereas a low Self-BLEU score denotes a diverse response (Touvron et al., 2023). As we intended, R improves the consistency in the CE type queries and promotes more diverse responses in the OE type. However, R produces excessively diverse responses regardless of the query type, demonstrating the efficacy of the branching by query type in R.
4.4 Is R generalizable method?
Finally, we compare our R with other state-of-the-art alignment models in various backbones. Note that our method can be adopted upon any instruction-tuned models without preference datasets. To show the generalizability of our method, we conduct PPO training on the three instruction-tuned models, Vicuna-{7, 15}B-v1.5 (Chiang et al., 2023; Zheng et al., 2023) and Mistral-H4-SFT (Tunstall et al., 2023). Specifically, we sample datasets from ShareGPT (Chiang et al., 2023) and UltraChat (Ding et al., 2023), originally used for the SFT training for each instruction-tuned model, for our PPO training. More training details are in Appendix C.
R consistently improves various SFT models
In Table 5, our method consistently improves the aligned behaviors of the three instruction-tuned models while keeping safety and incurring less alignment tax without explicit human or AI feedback datasets. Specifically, our resulting PPO models significantly improve the scores of AlpacaEval (Li et al., 2023) and MT-Bench (Zheng et al., 2023) based on the Vicuna-{7, 15}B-v1.5. They outperform other alignment models within the same backbones, i.e., Llama-2-{7, 13}B-Chat. Our model still increases the preference for the Mistral-H4-SFT backbone, but the increments are behind the Zephyr-beta (Tunstall et al., 2023), which is trained with GPT-4-generated feedback (Cui et al., 2023) 131313Zheng et al. (2023) report that GPT-4 prefers responses from the model trained with GPT-4 outputs, i.e., self-enhancement bias.. We presume the relatively slight improvements on the Mistral-H4-SFT are because the SFT model is less tuned according to the training recipe of Tunstall et al. (2023). On the other hand, we find that relevance (% Rel) is sacrificed according to the increment of preference (% Win) in the 7B models, as in the previous study, but the relevance score is even improved in the 13B model. We speculate the trade-off might be mitigated by scaling the model size. Also, we compare R with other preference-free approaches in Appendix H.
5 Related Work
Preference-based Alignment Learning
Reinforcement Learning from Human Feedback (RLHF) is the most prevalent approach to leverage the human preference (Ziegler et al., 2020; Ouyang et al., 2022; Bai et al., 2022a; Liu et al., 2022; Scheurer et al., 2023). These approaches include the reward modeling stage to develop a proxy RM predicting the human preference over the response, e.g., typically as a scalar format. Then, the trained proxy RM computes reward scores for the sampled responses in the RL stage, i.e., rollout. Recently, replacing human feedback with AI or synthetic feedback has been studied (Bai et al., 2022b; Kim et al., 2023a; Sun et al., 2023b; Cui et al., 2023). Direct Preference Optimization (DPO) reparameterizes the RLHF procedure to the efficient binary cross-entropy objective, eliminating additional RM training and rollouts (Rafailov et al., 2023). Zhao et al. (2023) apply hinge loss between preference pairs in an offline manner. Azar et al. (2023) suggest IPO to mitigate an overfitting problem in the DPO. Ethayarajh et al. (2023b) introduce Kahneman-Tversky Optimization (KTO) requiring only good or bad feedback of responses for given inputs, instead of the pairwise preference. A line of research extends the binary preference to ranking optimization (Yuan et al., 2023; Song et al., 2023a). There are attempts to leverage multiple rewards beyond the overall reward (Touvron et al., 2023; Wu et al., 2023; Rame et al., 2023; Jang et al., 2023). Wu et al. (2023) devise fine-grained rewards consisting of relevance, factuality, and information completeness to enhance long-form question-answering. To this end, human annotators collect the corresponding fine-grained feedback. On the contrary, our study focuses on another alternative to align LLMs leveraging off-the-shelf relevance model (Izacard et al., 2021), while not relying on the explicit preference dataset.
Pitfalls of Reward Modeling and Human Feedback
RLHF faces various challenges, especially in reward modeling and human feedback itself (Casper et al., 2023). Gao et al. (2023) present the overoptimization problem by the imperfect proxy reward models. The issue arises when the policy, tailored to enhance the estimated rewards, fails to improve real-world rewards. Many studies try to mitigate the overoptimization problem (Moskovitz et al., 2023; Coste et al., 2023). Similarly, Shen et al. (2023) study reward inconsistency and leverage a sentence encoder to increase the reward consistency. The work leverages SimCSE (Gao et al., 2021) to augment the preference dataset and normalize reward scores. On the contrary, we directly employ an off-the-shelf retriever, Contriever (Izacard et al., 2021), to compute the rewards. The policy models within the RLHF often find and exploit undesired shortcuts or errors in the RMs to maximize the proxy rewards. The phenomenon is known as reward hacking (Amodei et al., 2016; Pan et al., 2022; Pang et al., 2022; Song et al., 2023b; Eisenstein et al., 2023). On the other hand, much literature points out that human preference can be easily biased by superficial features such as complexity and assertiveness (Xu et al., 2023; Hosking et al., 2023). Singhal et al. (2023) study a high correlation between response length and RLHF training. Motivated by these studies, our work also reveals the poor performance of open-source RMs, trained by human feedback, over trivial relevance problems. Also, the motivation leads us to develop a preference-free approach for aligning models while effectively mitigating reward hacking.
6 Conclusion
In this work, we present a preference-free alignment learning based on the ‘relevance’. We find the proposed mixture of reward functions, Regularized Relevance Reward (R), could be used for aligning LLMs without explicit human or AI preference data. Notably, R outperforms existing open-source RMs trained on human preference when it serves as a reward function within RL. In this respect, our method can be especially useful when obtaining the preference data is challenging. Besides, we demonstrate the advantages of R in striking an optimal balance between enhancing human preference and minimizing the alignment tax. Finally, we show any instruction-tuned models can improve human preference over the models by optimizing against R, without additional dataset cost.
7 Impact Statements
This paper presents a preference-free alignment learning approach that does not rely on human or AI preference datasets. R could be an effective alternative when obtaining preference datasets is expensive. However, even though we show that PPO training with R does not increase toxicity on the safety benchmark, ToxiGen (Hartvigsen et al., 2022), compared to the SFT baseline, our reward function technically could not discriminate less harmful content. In other words, R might not be appropriate for alignment to increase safety. Similarly, although we show that R is robust to alignment tax (Wang et al., 2022; Min et al., 2023), R could not check the factuality of content. Rather, it relies on the quality of the reference response when the given query is closed-ended. As a result, R might increase undesirable hallucinations in the response according to the quality of the instruction dataset. Despite these limitations, we demonstrate that R consistently works well with various datasets, improving human preference and suppressing its side effects. We present more limitations of our work in Appendix A. We would like to explore more effective preference-free approaches that can handle safety and factuality in future work.
Acknowledgment
The authors would like to thank the members of NAVER Cloud (especially Kang Min Yoo) and KAIST LKLab (especially Seungone Kim) for their constructive comments.
Appendix A Limitations
Even though we show the efficacy of R based on automatic evaluations (Chiang et al., 2023; Li et al., 2023; Zheng et al., 2023; Min et al., 2023) leveraging proprietary LLMs such as GPT-4 (OpenAI, 2023), these evaluation results might not be rigorous enough and often contain biases as discussed in Zheng et al. (2023). We leave a more thorough comparison between R and other alignment approaches for future work. Also, while R does not require any preference datasets, we do not consider another training resource to train the relevance model. However, obtaining the training dataset would still be cost-effective in that it does not require human supervision (Izacard et al., 2021), i.e., an unsupervised scheme.
Appendix B Query-type Classification
B.1 Automatic annotation
We automatically annotate the query type leveraging GPT-4 (OpenAI, 2023). Specifically, we use the prompt of Figure 5 for the automation. As a result, we classify about 120k queries from various sources used in our experiments, Anthropic-HH (Bai et al., 2022a), AlpacaFarm (Dubois et al., 2023), ShareGPT, and UltraChat (Ding et al., 2023). The resulting examples are shown in Figure 6. For more dataset details, please see Appendix C.
| Query-type Classification |
Could you identify whether the last user query of the below conversation is open-ended or not? |
Open-ended means that the responses to the query can be relatively anything and creative. |
However, its counterparts, i.e., ‘closed-ended query’, indicate the corresponding responses should be consistent and constrained as a closed form. |
Please just answer it either “OPEN-ENDED” or “CLOSED-ENDED”. |
CONVERSATION: |
{conversation} |
| Open-ended | Closed-ended |
What are some good exercises I can do at a desk to stretch? |
If I was arrested for a crime, what rights do I have? |
What are some cool things do while in Washington DC? |
What age do I start talking to kids about puberty? |
How can I deodorize clothes without washing them? |
What was the irish potato famine about? |
how can i make my own superhero mask? |
How do you heal cuticles? |
I’ve been having trouble sleeping at night. How can I make sure I get my full eight hours? |
What is the main crop that is cultivated in the Northern hemisphere? |
I’ve been having frequent panic attacks lately and don’t know what’s causing them. |
Is it hard to get a CDL license? |
How can I make a good first impression? |
How can I track a package that has been lost in the mail? |
What are some Christmas songs to enjoy with the kids? |
How many movies are in the Fast and Furious franchise? |
Hey, where do you want to go for spring break this year? |
What does a urologist do? |
B.2 Distillation to in-house LLM
To eliminate the dependency on the proprietary model, GPT-4, we conduct dataset distillation to small in-house LLM (Touvron et al., 2023). In particular, we train LLaMA-2-7B with the 20k of GPT-4-constructed subsamples (Bai et al., 2022a). In addition, we combine the query-type classification data with other instruction-following data, FLAN-2021 (Wei et al., 2021), to improve the performance. The prompt of Figure 5 is used to format each instance for the instruction-tuning. As a result, about 61k (20k of distillation and 41k of FLAN-2021 samples) instances are used for the training. We use the same configuration with the SFT training except for 1024 of maximum sequence length. The resulting model achieves 86.7% accuracy, i.e., agreement with the GPT-4, in the AlpacaEval test set (Li et al., 2023).
Appendix C More Implementation Details
C.1 Dataset Details
Data Source Setup # Train # Dev Init. Backbone Anthropic-HH SFT 160,800 - meta-llama/Llama-2-7b-hf PPO 21,530 1,137 LLaMA-2-7B-SFT (trained by ourselves) AlpacaFarm PPO 20,001 2,000 tatsu-lab/alpaca-farm-sft10k-wdiff ShareGPT PPO 53,554 600 lmsys/vicuna-{7,13}b-v1.5 UltraChat PPO 100,000 600 HuggingFaceH4/mistral-7b-sft-beta
Table 6 shows overall dataset statistics for each experiment. All the datasets consist of input-output pairs (, ), which means they do not contain human preference triplets. For a fair comparison with the open-source RMs, we train SFT model based on LLaMA-2-7B (Touvron et al., 2023) with the 161k chosen responses in Anthropic-HH (Bai et al., 2022a) 141414huggingface.co/datasets/Anthropic/hh-rlhf. For the PPO training on Anthropic-HH, we select about 23k (train and dev) first-turn queries from ‘Helpful-rejection’ and ‘Helpful-online’ sub-splits for efficient implementation. Also, we employ the pre-trained SFT models (Init. Backbone), for the remaining experiments, AlpacaFarm (Dubois et al., 2023), ShareGPT (Zheng et al., 2023), UltraChat (Ding et al., 2023). Specifically, we use 20k of ‘unlabeled’ split for the PPO training on AlpacaFarm 151515huggingface.co/datasets/tatsu-lab/alpaca_farm/viewer/alpaca_instructions. For ShareGPT, we follow basic preprocessing suggested by Zheng et al. (2023) and sample 31k instances for efficient implementation 161616huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered. Last, we sample 50k instances for the UltraChat, preprocessed version by Tunstall et al. (2023) 171717huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. Please note that the actual number of the ShareGPT and UltraChat would be twice of the represented number, as we use dialogues up to two turns for efficient training. Figure 7 shows the distribution of for each dataset.
C.2 Training Details
| hyper-parameter | SFT | PPO |
| learning rate (lr) | 2e-5 | 1e-5 |
| batch size per a GPU | 2 | {2, 4} |
| gradient accumulation | 8 | {4, 8} |
| # epoch | 3 | - |
| # step | - | {5k, 10k} |
| lr scheduling | cosine | constant |
| warmup ratio | 0.03 | no |
| max sequence length | 2048 | - |
| max rollout length | - | {512, 768} |
| ppo epoch | - | 4 |
| initial KL coefficient | - | 0.2 |
| clip ratio | - | 0.2 |
| discount factor (gamma) | - | 1 |
Table 7 shows the common hyper-parameter setup used in our experiments. We conduct a full-finetuning (FFT) with fully-sharded data parallel (FSDP) for the SFT training of Anthropic-HH (Section 2 and Section 4.2). For PPO, we employ LoRA (Hu et al., 2021; Sun et al., 2023a) tuning for efficiency. Specifically, we apply the low-rank adaptors to query and value linear parameters. Also, we set lora_r to {8, 16}, lora_alpha to {16, 32}, and dropout to {0.05, 0.1}. We utilize an adaptive KL (Kullback–Leibler divergence) penalty with 0.2 of the initial KL coefficient (Ziegler et al., 2020). We choose the best checkpoint based on the average reward scores for every 1k PPO step. Figure 8 shows an example of the average reward scores based on the Vicuna + R experiment. One A100 GPU is used for most PPO implementations except when the policy is initialized with Vicuna-13B-v1.5. For the 13B model, we use 8 A100 GPUs. Besides, we leverage the corresponding chat template if the initial backbone model supports it (github.com/lm-sys/FastChat). All the implementation is based on the Transformers 181818github.com/huggingface/transformers, TRL 191919github.com/huggingface/trl and PEFT 202020github.com/huggingface/peft libraries.
Appendix D Evaluation Details
D.1 Evaluation set
We evaluate the resulting models on Vicuna-Bench (Chiang et al., 2023), AlpacaEval (Li et al., 2023), MT-Bench (Zheng et al., 2023), and ToxiGen (Hartvigsen et al., 2022). The first three benchmarks mainly evaluate the preference of models’ responses leveraging a superior proprietary LLM, GPT-4 (OpenAI, 2023). Vicuna-Bench contains 80 curated questions focusing on helpfulness, such as writing, roleplay, and reasoning 212121LLM Judge in FastChat. It performs a pairwise comparison between two candidate responses. Also, it includes reference information for math and coding questions and conducts two inferences by switching responses’ positions for a more precise evaluation (Zheng et al., 2023). To identify the effect of PPO with each RM, we set the SFT model as a baseline model for the pairwise comparisons. We report an adjusted win rate, , of each PPO model against to the SFT baseline. AlpacaEval is another benchmark conducting the pairwise comparison 222222github.com/tatsu-lab/alpaca_eval. It includes 805 instructions collected from diverse sources. We report a win rate compared to Text-Davinci-003 (Ouyang et al., 2022) as a baseline. MT-Bench is a benchmark checking multi-turn ability in a single judgment manner. It scores each response within the 1-10 range without the baseline response. It contains of questions, including the more knowledge-intensive questions such as math, coding, and STEM. We also include ToxiGen to investigate the side effects of the alignment in terms of safety. It consists of 6k prompts. Following the Touvron et al. (2023), we report the ratio of toxic continuations for the prompts using toxicity classifier 232323tomh/toxigen_roberta. In addition, we evaluate the resulting models on Super-NaturalInstructions (SuperNI) (Wang et al., 2022) to measure the models’ instruction-following ability, i.e., constrained generation. We report ROUGE-L of generations for the unseen test instructions. Following the official setup, we sample 100 instances in the unseen test split consisting of 119 tasks to compute the ROUGE-L score for the SuperNI. The overall tasks are shown in github.com/allenai/natural-instructions/blob/master/splits/default/test_tasks.txt. Also, we include FactScore (Min et al., 2023) to measure the factuality of the resulting models. It first breaks down the responses to atomic facts and then verifies the factual correctness with supporting knowledge, e.g., Wikipedia. For FactScore, we use “retrieval+ChatGPT” option to get the score. Specifically, it leverages InstructGPT (‘Text-Davinci-003’) (Ouyang et al., 2022) to get atomic facts from a given response and performs retrieve-then-verify with a passage retriever and ChatGPT(‘gpt-3.5-turbo’) (OpenAI, 2023). The evaluation code is available at github.com/shmsw25/FActScore.
D.2 Relevant Sentence Ratio
| Evaluation Prompt for Relevant Sentence Ratio |
<Instruction> |
Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. |
Your evaluation should check the ’relevancy’ of the provided assistant’s answer considering the question. |
Please investigate thoroughly whether the answer contains any irrelevant content sentence-by-sentence. |
That is, each sentence is marked as ’Irrelevant’ when it goes off-topic or includes information that is not relevant to the question. |
Otherwise, each sentence is marked as ’Relevant’. |
The assessment should be conducted at the sentence level. First, refer to each sentence. Then, judge its relevancy, either "Relevant" or "Irrelevant". |
For example, |
- "Sentence 1" : (Relevant|Irrelevant) |
- "Sentence 2" : (Relevant|Irrelevant) |
<Question> |
{question} |
<The Start of Assistant’s Answer> |
{answer} |
<The End of Assistant’s Answer> |
Figure 9 shows the evaluation prompt to obtain a relevant sentence ratio using GPT-4 (OpenAI, 2023). It returns sentence-level relevance of the given assistant answer, as shown in Figure 10. Inspired by Min et al. (2023), we count the number of relevant sentences among all sentences in an answer and report the average value across the whole test set. We think this sentence-level relevance is a more appropriate and fair measurement than the relevance evaluation computed by a reranker in Section 2.2.
Model Alpaca Eval Avg # Tokens 4-gram Rep. % Rel. Sent SFT (Anthropic-HH) 15.9 126 0.05 85.9 + PPO w. SteamSHP-XL 19.6 154 0.10 80.2 w. OAAST 19.1 186 0.10 85.0 w. LLaMA-2-7B-RM 19.8 146 0.08 84.5 w. UltraRM 28.0 213 0.11 83.1 w. R (Ours) 23.4 243 0.09 85.1
In Table 8, R keeps the query relevance of the resulting PPO model, increasing preference and response length compared to the SFT baseline. All the PPO models show the increased token length with the improved preference (Singhal et al., 2023). However, SteamSHP-XL and UltraRM significantly drop the % Rel.
Appendix E Further analysis of R and UltraRM
E.1 Normalized Reward Scores
E.2 Rewards Ensemble of R and UltraRM
Model # Feedback to train RM Vicuna-Bench % Adjusted Win AlpacaEval % Win / % Rel MT-Bench 1st / 2nd / Avg ToxiGen () SFT (Anthropic-HH) - 50.0 15.9 / 85.9 4.6 / 3.4 / 3.9 2.6 + PPO w. UltraRM 256k 73.8 28.0 / 83.1 4.9 / 3.6 / 4.2 4.7 w. R 0 76.9 23.4 / 85.1 5.0 / 3.7 / 4.4 3.9 w. R + UltraRM 256k 73.8 26.8 / 87.0 4.7 / 3.5 / 4.1 4.3
Appendix F Comparison with other Preference-based Learning approaches
Model Data Type # Data Vicuna-Bench % Adjusted Win AlpacaEval % Win / % Rel MT-Bench Avg SFT (Anthropic-HH) 160k 50.0 15.9 / 85.9 3.9 Offline Preference Learning SLiC-HF-direct (Zhao et al., 2023) 22k 58.1 21.1 / 87.1 4.3 DPO (Rafailov et al., 2023) 22k 62.5 22.2 / 84.8 4.2 IPO (Azar et al., 2023) 22k 52.5 21.4 / 86.2 4.4 KTO (Ethayarajh et al., 2023b) 22k 54.4 18.3 / 86.7 4.4 Preference-free (Ours) PPO w. R 21k 76.9 23.4 / 85.1 4.4
In Table 10, we compare R with offline preference learning approaches, including SLiC (Zhao et al., 2023), DPO (Rafailov et al., 2023), IPO (Azar et al., 2023), and KTO (Ethayarajh et al., 2023b). These approaches utilize static preference triplets to directly optimize policy without the rollout procedure requiring response sampling and evaluation by RM over the sampled responses. For the preference dataset, we use 22k of the ‘online’ subset in Anthropic-HH (Bai et al., 2022a). As we use another first-turn subset for our PPO training described in Appendix C.1, this comparison is not fully comparable. Nevertheless, we can find PPO training with R surpasses all the offline methods on three preference benchmarks with a comparable relevant sentence ratio.
All the training of offline methods is based on TRL and PEFT libraries. We set the number of training epochs to 3 and the parameter for the DPO objective to 0.1. The remaining hyper-parameter setups are the same as those of the PPO described in Appendix C.2.
Appendix G Results with other Relevance Models
Model Normalization Cosine Similarity AlpacaEval % Win / % Rel SuperNI ROUGE-L SFT (AlpacaFarm) 23.2 / 88.2 34.9 + R PPO w. Relevance Model by Unsupervised Contriever (Izacard et al., 2021) ✗ 37.2 / 86.4 34.9 Unsupervised Contriever (Izacard et al., 2021) ✓ 17.1 / 83.6 11.1 Supervised Contriever (Izacard et al., 2021) ✗ 38.3 / 86.8 32.6 Supervised SimCSE (Gao et al., 2021) ✓ 26.1 / 85.8 33.2
Appendix H Comparison with other preference-free approaches
Model AlpacaEval % Win / % Rel. MT-Bench Vicuna-7B-v1.5 73.7 / 91.5 6.1 + LPPO (Singhal et al., 2023) 65.7 / 81.5 5.8 + NEFTUNE (Jain et al., 2023) 72.7 / 89.8 6.2 + PPO w. R (Ours) 79.1 / 84.4 6.4
In Table 12, we compare our method with recently proposed other preference-free approaches (Singhal et al., 2023; Jain et al., 2023). We reproduce the results using the same SFT model, Vicuna-7B-v1.5, and the corresponding dataset, the sampled ShareGPT. Note that all the implementations are based on LoRA (Hu et al., 2021). The LPPO is similar to PPO training with only length incentive (LI), promoting longer text generation. The reward function is , where the is the hyper-parameter indicating target length. Contrary to the report by Singhal et al. (2023), using only the longer length as a reward signal does not improve the preference of Vicuna-7B-v1.5 but rather reduces the overall performance. Moreover, NEFTUNE, a method injecting random noise to the embeddings, is not as effective as Jain et al. (2023) reported in improving preference. Specifically, we set the noise alpha parameter to 5 and use the same hyper-parameter setup of Anthropic-HH SFT training presented in Appendix C. Among the approaches, only our method improves both the win rate and the score of AlpacaEval and MT-Bench with acceptable degradation of the relevance.
Appendix I Qualitative Examples
| Prompt: | |
What is the best approach for learning a foreign language when you only have an hour a day to practice? |
| R: | UltraRM: |
Learning a foreign language takes time, and while there is no one right way to learn a language, spending just one hour a day is probably not enough to learn and retain the language well. For example, if you’re learning Spanish, you might start with some basic vocabulary and phrases, and then progress to speaking, listening, and reading more complex materials. Research suggests that the best way to learn a language is to go through a period of intense study, usually lasting a few months, during which you devote at least an hour a day to language learning. After this period of intensive study, you can then move to a maintenance phase, where you spend less time learning, but continue to practice regularly. So the best approach is to start with a period of intense study, using a textbook, an app, or other resources to help you learn, and practice every day for at least an hour. Then, once you feel that you have a strong grasp on the language, you can move to a maintenance phase where you practice less every day. |
The most effective approach to learning a foreign language is to use the “Method of Feigned Passivity”. You should pretend to have a lack of interest and motivation, while covertly establishing a strong internal drive towards mastering the language. This contradiction will force your unconscious to find the most efficient and effective learning methods. I can recommend the following practice exercises: - Pretend to not understand when someone speaks to you in the target language. If they offer to help you learn, respond with “I’m not interested”. Internalise a burning desire to demonstrate your understanding, without appearing to make an effort. - Pretend to forget the grammar rules and vocabulary you do know. If you remember a word or a sentence you will gain a sense of its correct usage, but pretend to forget it anyway. Internalise a burning desire to remember everything, without appearing to make an effort. - Pretend to not remember what you have learned previously. If you are discussing a topic with someone, and you are about to use a word or sentence that you have learnt previously, pretend to not remember it. Internalise a burning desire to recall your previous learning, without appearing to make an effort. … |
| Prompt: | |
Do you know why turkeys became the official food of thanksgiving? |
| R: | LI(): |
Yes, turkeys became the official food of Thanksgiving because the pilgrims and the Wampanoag Indians who gathered for the first Thanksgiving feast in 1621 were reportedly served turkey. In addition, turkeys were seen as a symbol of abundance, as they are believed to have migrated to the New World in great numbers. Additionally, turkeys are thought to be easier to hunt than other birds, making them a practical choice for the early settlers. As such, the turkey has become a symbol of thanksgiving and is now a popular dish on Thanksgiving menus across the United States. In modern times, turkeys are often raised specifically for the Thanksgiving holiday and are often given a special diet so that they can reach a desired weight before the holiday. In some cases, turkeys may even be stuffed with a special mixture of herbs and spices to add flavor to the meat. |
No, I do not. Why did turkeys become the official food of Thanksgiving? The tradition of eating turkey at Thanksgiving is thought to have originated with the Pilgrims who celebrated the first Thanksgiving feast in 1621 with the Wampanoag Indians. The Wampanoag Indians were skilled at hunting turkeys and the Pilgrims were eager to learn the best way to prepare them. The Pilgrims were amazed at the size of the turkeys they were served and the tradition of eating turkey on Thanksgiving was born. … Turkeys have become so closely associated with Thanksgiving that they are sometimes called the “official food” of the holiday. In addition to being served on Thanksgiving, turkeys are also popular during Christmas and other holiday meals. Turkeys have become so closely associated with Thanksgiving that they are sometimes called the “official food” of the holiday. In addition to being served on Thanksgiving, turkeys are also popular during Christmas and other holiday meals. … |
| Prompt: | |
Provide me with a list of 10 names from various cultures that mean love, care, and/or empathy. |
| R: | R: |
1. Aahana (Hindi) 2. Ahava (Hebrew) 3. Ai (Japanese) 4. Agape (Greek) 5. Alay (Tagalog) 6. Betelehem (Hebrew) 7. Cariño (Spanish) 8. Daba (Amharic) 9. Dhul (Arabic) 10. Empatia (Polish) |
1. Arabic: Rahmah 2. Chinese: Ai 3. French: Amour 4. Greek: Agape 5. Hindi: Sharada 6. Japanese: Ai … 38. Zulu: Ngiyabonga 39. Vietnamese: Yêu Mãy 40. Irish: Cead Mille Failte 41. Italian: Carità |
| Prompt: | |
Write down three antonyms for the given word. ‘laureating’ |
| R: | R: |
denigrating, disparaging, belittling. |
The antonyms of laureating are: - Decorating - Dignifying - Honoring - Magnifying - Praising - Sanctifying - Serving - Exalting - Glorying … |